John received a parcel of 2 dozen mangoes.six of them were rotten.What fraction of the mangoes was good?

Grade:6

1 Answers

Rohit Bamel
6 years ago
Total Mangoes =24Rotten Mangoes = 24-6 = 18Fraction of Rotten mangoes = 6/24*100 =25%Now ,Fraction of good Mangoes = 100-25 =75%